# thanks: Corresponding Author

### v1

@@ III.1 $\mathfrak{M\rightarrow}J/\psi B_{c}^{+}$
$\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})=\int_{(3m_{c}+m_{b})^{2}}^{s_{0}}ds\int_{(m_{b}+m_{c})^{2}}^{s_{0}^{\prime}}ds^{\prime}$
$\times\rho_{1}(s,s^{\prime},q^{2})e^{-s/M_{1}^{2}}e^{-s^{\prime}/M_{2}^{2}},$ (22)
@@
The function $\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})$ depends on two pairs of parameters $(M_{1}^{2},s_{0})$ and $(M_{2}^{2},s_{0}^{\prime})$. The first of them corresponds to the channel of the molecule $\mathfrak{M}$, while $(M_{2}^{2},s_{0}^{\prime})$ is related to the channel of $B_{c}^{+}$ meson. They are restricted in accord with standard rules of SR calculations. Our studies prove that Eq. (11) for the parameters $(M_{1}^{2},s_{0})$ and $(M_{2}^{2},s_{0}^{\prime})$
@@
It is known that the SR method leads to credible results in the Euclidean region $q^{2}<0$, whereas $g_{1}(q^{2})$ amounts $g_{1}$ at $q^{2}=m_{J/\psi}^{2}$. Hence, it is convenient to make use of new function $g_{1}(Q^{2})$ where $Q^{2}=-q^{2}$, and employ it for subsequent investigations. The QCD results for $g_{1}(Q^{2})$ where $Q^{2}=2-30~\mathrm{GeV}^{2}$are plotted in Fig. 3.
@@
$g_{1}\equiv\mathcal{Z}_{1}(-m_{J/\psi}^{2})=(2.6\pm 0.5)\times 10^{-1}\ \mathrm{GeV}^{-1}.$ (26)
@@
$|M_{1}|^{2}$ $=$ $\frac{1}{4m_{B_{c}}^{2}}\left[m^{6}-2m^{4}m_{J/\psi}^{2}+2m_{B_{c}}^{2}(m_{B_{c}}^{2}-m_{J/\psi}^{2})^{2}\right$. (28)
$\left.+m^{2}(m_{J/\psi}^{4}+6m_{J/\psi}^{2}m_{B_{c}}^{2}-3m_{B_{c}}^{4})\right].$
@@
Figure 3: SR data and fit functions $\mathcal{Z}_{1}(Q^{2})$ (solid curve) and $\mathcal{Z}_{2}(Q^{2})$ (dashed line). The triangle and circle denote the points $Q^{2}=-m_{J/\psi}^{2}$ and $Q^{2}=-m_{\eta_{c}}^{2}$, respectively.
@@ III.2 $\mathfrak{M\rightarrow}\eta_{c}B_{c}^{\ast+}$
$\Pi_{\mu\nu}^{2}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{I_{\mu}^{B_{c}^{\ast}}(y)$ (31)
$\times I^{\eta_{c}}(0)I_{\nu}^{\dagger}(x)\}|0\rangle.$
@@
$\Pi_{\mu\nu}^{2\mathrm{Phys}}(p,p^{\prime})=\frac{g_{2}(q^{2})\Lambda f_{\eta_{c}}m_{\eta_{c}}^{2}f_{B_{c}^{\ast}}m_{B_{c}^{\ast}}}{2m_{c}\left(p^{2}-m^{2}\right)\left(p^{\prime 2}-m_{B_{c}^{\ast}}^{2}\right)(q^{2}-m_{\eta_{c}}^{2})}$
$\times\left[\frac{m^{2}-m_{B_{c}^{\ast}}^{2}+q^{2}}{2}g_{\mu\nu}-p_{\mu}p_{\nu}+p_{\mu}^{\prime}p_{\nu}\right$.
$\left.+\frac{m^{2}+m_{B_{c}^{\ast}}^{2}-q^{2}}{m_{B_{c}^{\ast}}^{2}}p_{\mu}p_{\nu}^{\prime}\right]+\cdots.$ (33)
@@
$\langle 0|I_{\mu}^{B_{c}^{\ast}}|B_{c}^{\ast+}(p^{\prime},\varepsilon)\rangle$ $=$ $f_{B_{c}^{\ast}}m_{B_{c}^{\ast}}\varepsilon_{\mu},$
$\langle 0|I^{\eta_{b}}|\eta_{c}(q)\rangle$ $=$ $\frac{f_{\eta_{c}}m_{\eta_{c}}^{2}}{2m_{c}},$ (34)
@@
$\langle\eta_{c}(q)B_{c}^{\ast+}(p^{\prime},\varepsilon)|\mathfrak{M}(p,\epsilon)\rangle=g_{2}(q^{2})\left[(p\cdot q)(\epsilon\cdot\varepsilon^{\ast})\right$.
$\left.-(q\cdot\epsilon)(p\cdot\varepsilon^{\ast})\right],$ (35)
@@
$\Pi_{\mu\nu}^{2\mathrm{OPE}}(p,p^{\prime})=\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\mathrm{Tr}\left[\gamma_{\mu}S_{c}^{ia}(y-x)\right$.
$\left.\times\gamma_{\nu}S_{c}^{aj}(x)\gamma_{5}S_{c}^{jb}(-x)\gamma_{5}S_{b}^{bi}(x-y)\right].$ (36)
@@
$g_{2}(q^{2})=\frac{4m_{c}}{\Lambda f_{\eta_{c}}m_{\eta_{c}}^{2}f_{B_{c}^{\ast}}m_{B_{c}^{\ast}}}\frac{q^{2}-m_{\eta_{c}}^{2}}{m^{2}-m_{B_{c}^{\ast}}^{2}+q^{2}}$
$\times e^{m^{2}/M_{1}^{2}}e^{m_{B_{c}^{\ast}}^{2}/M_{2}^{2}}\Pi_{2}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}),$ (37)
@@ IV.1 $\mathfrak{M\rightarrow}B^{\ast+}D^{0}$ and $B^{\ast 0}D^{+}$
Here, we study investigate the mode $\mathfrak{M\rightarrow}B^{\ast+}D^{0}$ and calculate its partial decay width. We plan to determine the strong coupling $g_{3}$ of particles at $\mathfrak{M}B^{\ast+}D^{0}$. The correlation function to be considered is
$\Pi_{\mu\nu}^{3}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{I_{\mu}^{B^{\ast}}(y)$ (42)
$\times I^{D^{0}}(0)I_{\nu}^{\dagger}(x)\}|0\rangle,$
@@
$\langle 0|I_{\mu}^{B^{\ast}}|B^{\ast+}(p^{\prime},\varepsilon)\rangle$ $=$ $f_{B^{\ast}}m_{B^{\ast}}\varepsilon_{\mu},$
$\langle 0|I^{D^{0}}|D^{0}(q)\rangle$ $=$ $\frac{f_{D}m_{D^{0}}^{2}}{m_{c}}.$ (44)
@@
$\Pi_{\mu\nu}^{3\mathrm{OPE}}(p,p^{\prime})=\frac{1}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle\overline{c}c\rangle\mathrm{Tr}\left[\gamma_{\mu}S_{u}^{ij}(y)\right$.
$\left.\times\gamma_{5}S_{c}^{ja}(-x)\gamma_{\nu}\gamma_{5}S_{b}^{ia}(x-y)\right].$ (45)
Later on, we apply the relation
$m_{c}\langle\overline{c}c\rangle\approx-\frac{1}{12}\langle\frac{\alpha_{s}G^{2}}{\pi}\rangle$ (46)
extracted in the context of SR method Shifman:1978bx , and valid if terms in higher powers of $m_{c}^{-1}$ are small.
The form factor $g_{3}(Q^{2})$ is evaluated in the domain $Q^{2}=2-30\ \mathrm{GeV}^{2}$: At $Q^{2}=-m_{D^{0}}^{2}$ it equals to $g_{3}$. For $(M_{1}^{2},s_{0})$ we have used Eq. (11), and $(M_{2}^{2},s_{0}^{\prime})$ have changed in the regions
@@
$g_{3}\equiv\mathcal{Z}_{3}(-m_{D^{0}}^{2})=(6.3\pm 1.2)\times 10^{-2}~\mathrm{GeV}^{-1}$
.For the width of the process $\mathfrak{M\rightarrow}B^{\ast+}D^{0}$ our analysis predicts
@@ IV.2 $\mathfrak{M\rightarrow}$ $B^{+}D^{\ast 0}$ and $B^{0}D^{\ast+}$
The pair of the decay modes $\mathfrak{M\rightarrow}$ $B^{+}D^{\ast 0}$ and $\mathfrak{M\rightarrow}$ $B^{0}D^{\ast+}$ is explored in this subsection using the same techniques. Below, we give parameters of the process $\mathfrak{M\rightarrow}$ $B^{+}D^{\ast 0}$ bearing in mind that the second mode approximately shares them amounts to parameters of the first channel.
@@
$\Pi_{\mu\nu}^{4}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{I^{B^{+}}(y)$ (50)
$\times I_{\mu}^{D^{\ast 0}}(0)I_{\nu}^{\dagger}(x)\}|0\rangle.$
@@
$\langle 0|I^{B^{+}}|B^{-}(p^{\prime})\rangle$ $=$ $\frac{f_{B}m_{B}^{2}}{m_{b}},$
$\langle 0|I_{\mu}^{D^{\ast 0}}|D^{\ast 0}(q,\varepsilon)\rangle$ $=$ $f_{D^{\ast 0}}m_{D^{\ast 0}}\varepsilon_{\mu},$ (52)
@@
$\langle B^{+}(p^{\prime})D^{\ast 0}(q,\varepsilon)|\mathfrak{M}(p,\epsilon)\rangle=g_{4}(q^{2})\left[(p\cdot p^{\prime})(\epsilon\cdot\varepsilon^{\ast})\right$.
$\left.-(p^{\prime}\cdot\epsilon)(p\cdot\varepsilon^{\ast})\right].$ (53)
@@
$\Pi_{\mu\nu}^{4\mathrm{OPE}}(p,p^{\prime})=\frac{1}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle\overline{c}c\rangle\mathrm{Tr}\left[\gamma_{5}S_{u}^{ij}(y)\right$.
$\left.\times\gamma_{\mu}S_{c}^{ja}(-x)\gamma_{\nu}\gamma_{5}S_{b}^{ai}(x-y)\right].$ (54)
Numerical computations of $g_{4}(Q^{2})$ is performed at $Q^{2}=2-30\ \mathrm{GeV}^{2}$ using the regions Eq. (11) for $(M_{1}^{2},s_{0})$ and
@@
for the Borel and continuum subtraction parameters $(M_{2}^{2},s_{0}^{\prime})$. The extrapolating function $\mathcal{Z}_{4}(Q^{2})$ that has been employed to compute the coupling $g_{4}(Q^{2})$ is fixed by the following constants $\mathcal{Z}_{4}^{0}=0.060~\mathrm{GeV}^{-1},z_{4}^{1}=3.57,z_{4}^{2}=-2.50.$ This function leads to the predictions

### v2

@@ III.1 $\mathfrak{M\rightarrow}J/\psi B_{c}^{+}$
$\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})=\int_{(3m_{c}+m_{b})^{2}}^{s_{0}}ds\int_{(m_{b}+m_{c})^{2}}^{s_{0}^{\prime}}ds^{\prime}\rho_{1}(s,s^{\prime},q^{2})e^{-s/M_{1}^{2}}e^{-s^{\prime}/M_{2}^{2}},$ (22)
@@
The function $\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})$ depends on two pairs of parameters $(M_{1}^{2},s_{0})$ and $(M_{2}^{2},s_{0}^{\prime})$. The first of them corresponds to the channel of the molecule $\mathfrak{M}$, while $(M_{2}^{2},s_{0}^{\prime})$ is related to the channel of $B_{c}^{+}$ meson. They are restricted in accord with standard rules of SR calculations. Our studies prove that Eq. (11) for the parameters $(M_{1}^{2},s_{0})$ and
@@
It is known that the SR method leads to credible results in the Euclidean region $q^{2}<0$, whereas $g_{1}(q^{2})$ amounts to $g_{1}$ at $q^{2}=m_{J/\psi}^{2}$. Hence, it is convenient to make use of new function $g_{1}(Q^{2})$ where $Q^{2}=-q^{2}$, and employ it for subsequent investigations. The QCD results for $g_{1}(Q^{2})$ where $Q^{2}=2-30~\mathrm{GeV}^{2}$are plotted in Fig. 3.
@@
$g_{1}\equiv\mathcal{Z}_{1}(-m_{J/\psi}^{2})=(2.61\pm 0.50)\times 10^{-1}\ \mathrm{GeV}^{-1}.$ (26)
This result has been found by using Eq. (24). At the same time, SR data can be extended into region the region $Q^{2}<0$ by employing different extrapolating functions. To clarify this problem, we examine the fit function
$\mathcal{Z}_{1A}(Q^{2})=\frac{c_{0}}{\left(1-Q^{2}/m^{2}\right)^{4}[1-c_{1}(Q^{2}/m^{2})+c_{2}\left(Q^{4}/m^{4}\right)]},$ (27)
where $c_{0}$, $c_{1}$, and $c_{2}$ are fitted constants. We get $c_{0}=0.826\ \mathrm{GeV}^{-1}$, $c_{1}=1.844$, and $c_{2}=2.731$ from comparison of Eq. (27) and SR data. This function is presented in Fig. 3 as well, in which one sees that it agrees with SR data and $\mathcal{Z}_{1}(Q^{2})$. The fit function $\mathcal{Z}_{1A}(Q^{2})$ leads to the result $g_{1}=0.158~\mathrm{GeV}^{-1}$, which deviates $|0.003|$ from the value in Eq. (26). But this deviation is more than an order of magnitude smaller than uncertainties $\pm 0.05$ of $g_{1}$ arising from the sum rule method. For this reason, we neglect uncertainties appearing due to the choice of extrapolating functions and employ Eq. (24) in our computations. It is worth noting that analytical forms of the extrapolating functions $\mathcal{Z}_{1}(Q^{2})$ and $\mathcal{Z}_{1A}(Q^{2})$ are inspired by the SR for $g_{1}(q^{2})$ in Eq. (21), which contains both the exponential and $(m^{2}-q^{2})$-type factors.
@@
$|M_{1}|^{2}=\frac{1}{4m_{B_{c}}^{2}}\left[m^{6}-2m^{4}m_{J/\psi}^{2}+2m_{B_{c}}^{2}(m_{B_{c}}^{2}-m_{J/\psi}^{2})^{2}+m^{2}(m_{J/\psi}^{4}+6m_{J/\psi}^{2}m_{B_{c}}^{2}-3m_{B_{c}}^{4})\right].$ (29)
@@
Figure 3: Left panel: SR data and fit functions $\mathcal{Z}_{1}(Q^{2})$ (solid curve), $\mathcal{Z}_{1A}(Q^{2})$ (dashed line) and $\mathcal{Z}_{2}(Q^{2})$ (dot-dashed line). The triangle and circle denote values of the extrapolating functions at the points $Q^{2}=-m_{J/\psi}^{2}$ and $Q^{2}=-m_{\eta_{c}}^{2}$, respectively. Right panel: QCD data and extrapolating functions $\mathcal{Z}_{3}(Q^{2})$ (solid) and $\mathcal{Z}_{4}(Q^{2})$ (dashed). The rectangle and triangle fix fit functions at the points $Q^{2}=-m_{D^{0}}^{2}$ and $Q^{2}=-m_{D^{\ast 0}}^{2}$.
@@ III.2 $\mathfrak{M\rightarrow}\eta_{c}B_{c}^{\ast+}$
$\Pi_{\mu\nu}^{2}(p,p^{\prime})=i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{I_{\mu}^{B_{c}^{\ast}}(y)I^{\eta_{c}}(0)I_{\nu}^{\dagger}(x)\}|0\rangle.$ (32)
@@
$\Pi_{\mu\nu}^{2\mathrm{Phys}}(p,p^{\prime})$ $=$ $\frac{g_{2}(q^{2})\Lambda f_{\eta_{c}}m_{\eta_{c}}^{2}f_{B_{c}^{\ast}}m_{B_{c}^{\ast}}}{2m_{c}\left(p^{2}-m^{2}\right)\left(p^{\prime 2}-m_{B_{c}^{\ast}}^{2}\right)(q^{2}-m_{\eta_{c}}^{2})}\left[\frac{m^{2}-m_{B_{c}^{\ast}}^{2}+q^{2}}{2}g_{\mu\nu}-p_{\mu}p_{\nu}+p_{\mu}^{\prime}p_{\nu}\right$. (34)
$\left.+\frac{m^{2}+m_{B_{c}^{\ast}}^{2}-q^{2}}{m_{B_{c}^{\ast}}^{2}}p_{\mu}p_{\nu}^{\prime}\right]+\cdots.$
@@
$\langle 0|I_{\mu}^{B_{c}^{\ast}}|B_{c}^{\ast+}(p^{\prime},\varepsilon)\rangle=f_{B_{c}^{\ast}}m_{B_{c}^{\ast}}\varepsilon_{\mu},\ \ \langle 0|I^{\eta_{b}}|\eta_{c}(q)\rangle=\frac{f_{\eta_{c}}m_{\eta_{c}}^{2}}{2m_{c}},$ (35)
@@
$\langle\eta_{c}(q)B_{c}^{\ast+}(p^{\prime},\varepsilon)|\mathfrak{M}(p,\epsilon)\rangle=g_{2}(q^{2})\left[(p\cdot q)(\epsilon\cdot\varepsilon^{\ast})-(q\cdot\epsilon)(p\cdot\varepsilon^{\ast})\right],$ (36)
@@
$\Pi_{\mu\nu}^{2\mathrm{OPE}}(p,p^{\prime})=\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\mathrm{Tr}\left[\gamma_{\mu}S_{c}^{ia}(y-x)\gamma_{\nu}S_{c}^{aj}(x)\gamma_{5}S_{c}^{jb}(-x)\gamma_{5}S_{b}^{bi}(x-y)\right].$ (37)
@@
$g_{2}(q^{2})=\frac{4m_{c}}{\Lambda f_{\eta_{c}}m_{\eta_{c}}^{2}f_{B_{c}^{\ast}}m_{B_{c}^{\ast}}}\frac{q^{2}-m_{\eta_{c}}^{2}}{m^{2}-m_{B_{c}^{\ast}}^{2}+q^{2}}e^{m^{2}/M_{1}^{2}}e^{m_{B_{c}^{\ast}}^{2}/M_{2}^{2}}\Pi_{2}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}),$ (38)
@@ IV.1 $\mathfrak{M\rightarrow}B^{\ast+}D^{0}$ and $B^{\ast 0}D^{+}$
Here, we study the mode $\mathfrak{M\rightarrow}B^{\ast+}D^{0}$ and calculate its partial decay width. We plan to determine the strong coupling $g_{3}$ of particles at $\mathfrak{M}B^{\ast+}D^{0}$. The correlation function to be considered is
$\Pi_{\mu\nu}^{3}(p,p^{\prime})=i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{I_{\mu}^{B^{\ast}}(y)I^{D^{0}}(0)I_{\nu}^{\dagger}(x)\}|0\rangle,$ (43)
@@
$\langle 0|I_{\mu}^{B^{\ast}}|B^{\ast+}(p^{\prime},\varepsilon)\rangle=f_{B^{\ast}}m_{B^{\ast}}\varepsilon_{\mu},\ \ \ \langle 0|I^{D^{0}}|D^{0}(q)\rangle=\frac{f_{D}m_{D^{0}}^{2}}{m_{c}}.$ (45)
@@
$\Pi_{\mu\nu}^{3\mathrm{OPE}}(p,p^{\prime})=\frac{1}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle\overline{c}c\rangle\mathrm{Tr}\left[\gamma_{\mu}S_{u}^{ij}(y)\gamma_{5}S_{c}^{ja}(-x)\gamma_{\nu}\gamma_{5}S_{b}^{ia}(x-y)\right].$ (46)
The correlation function $\Pi_{\mu\nu}^{3\mathrm{OPE}}(p,p^{\prime})$ depends on three quark propagators and vacuum condensate $\langle\overline{c}c\rangle$, and differs from standard correlators which contain four quark propagators. In our computations we contract heavy and light quark fields. The mesons $B^{\ast+}D^{0}$ contain only one $c$-quark field, as a result, free $\overline{c}c$ quarks in $\mathfrak{M}$ form a local vacuum condensate. In other words, $\langle\overline{c}c\rangle$ appears in formulas instead of the quark propagator and is considering on equal footing with them.
In what follows, we apply the relation between the heavy quark and gluon condensates to express $\Pi_{\mu\nu}^{3\mathrm{OPE}}(p,p^{\prime})$ using relevant parameters. The available formula for $\langle\overline{c}c\rangle$ reads Shifman:1978bx ; Generalis:1983hb ; Bagan:1985zp
$m_{c}\langle\overline{c}c\rangle=-\frac{1}{12}\langle\frac{\alpha_{s}G^{2}}{\pi}\rangle+\frac{1}{m_{c}^{2}}\langle\frac{\alpha_{s}G^{3}}{\pi}\rangle\left(-\frac{1}{48}+\frac{13}{720}\right)+\cdots.$ (47)
In our analysis we use only the first term in Eq. (47) extracted in Ref. Shifman:1978bx , because next ones are suppressed by additional powers of $m_{c}^{-1}$ and are small.
The form factor $g_{3}(Q^{2})$ is evaluated in the domain $Q^{2}=2-20\ \mathrm{GeV}^{2}$: At $Q^{2}=-m_{D^{0}}^{2}$ it equals to $g_{3}$. For $(M_{1}^{2},s_{0})$ we have used Eq. (11), and $(M_{2}^{2},s_{0}^{\prime})$ have changed in the regions
@@
$g_{3}\equiv\mathcal{Z}_{3}(-m_{D^{0}}^{2})=(6.3\pm 1.2)\times 10^{-2}~\mathrm{GeV}^{-1}.$ (49)
For the width of the process $\mathfrak{M\rightarrow}B^{\ast+}D^{0}$ our analysis predicts
@@ IV.2 $\mathfrak{M\rightarrow}$ $B^{+}D^{\ast 0}$ and $B^{0}D^{\ast+}$
The pair of the decay modes $\mathfrak{M\rightarrow}$ $B^{+}D^{\ast 0}$ and $\mathfrak{M\rightarrow}$ $B^{0}D^{\ast+}$ is explored in this subsection using the same techniques. Below, we give parameters of the process $\mathfrak{M\rightarrow}$ $B^{+}D^{\ast 0}$ bearing in mind that the second mode approximately shares those of the first channel.
@@
$\Pi_{\mu\nu}^{4}(p,p^{\prime})=i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{I^{B^{+}}(y)I_{\mu}^{D^{\ast 0}}(0)I_{\nu}^{\dagger}(x)\}|0\rangle.$ (52)
@@
$\langle 0|I^{B^{+}}|B^{-}(p^{\prime})\rangle=\frac{f_{B}m_{B}^{2}}{m_{b}},\ \ \langle 0|I_{\mu}^{D^{\ast 0}}|D^{\ast 0}(q,\varepsilon)\rangle=f_{D^{\ast 0}}m_{D^{\ast 0}}\varepsilon_{\mu},$ (54)
@@
$\langle B^{+}(p^{\prime})D^{\ast 0}(q,\varepsilon)|\mathfrak{M}(p,\epsilon)\rangle=g_{4}(q^{2})\left[(p\cdot p^{\prime})(\epsilon\cdot\varepsilon^{\ast})-(p^{\prime}\cdot\epsilon)(p\cdot\varepsilon^{\ast})\right].$ (55)
@@
$\Pi_{\mu\nu}^{4\mathrm{OPE}}(p,p^{\prime})=\frac{1}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle\overline{c}c\rangle\mathrm{Tr}\left[\gamma_{5}S_{u}^{ij}(y)\gamma_{\mu}S_{c}^{ja}(-x)\gamma_{\nu}\gamma_{5}S_{b}^{ai}(x-y)\right].$ (56)
Numerical computations of $g_{4}(Q^{2})$ is performed at $Q^{2}=2-20\ \mathrm{GeV}^{2}$ using the regions Eq. (11) for $(M_{1}^{2},s_{0})$ and
@@
for the Borel and continuum subtraction parameters $(M_{2}^{2},s_{0}^{\prime})$. The extrapolating function $\mathcal{Z}_{4}(Q^{2})$ that has been employed to compute the coupling $g_{4}(Q^{2})$ is fixed by the following constants $\mathcal{Z}_{4}^{0}=0.060~\mathrm{GeV}^{-1},z_{4}^{1}=3.57,z_{4}^{2}=-2.50$ and shown also in Fig. 3. This function leads to the predictions